# QCSE: A Pretrained Quantum Context-Sensitive Word Embedding for Natural Language Processing

### v1

@@ III-B2 Ansatz Circuit and Trainable Parameters
The second component of the QCSE model is the ansatz circuit, which comprises trainable parameters. The architecture of the ansatz circuit is shown in Figure 3 in the red block. Let $\mathbf{\Theta}=[\boldsymbol{\theta}_{1},\boldsymbol{\theta}_{2},\dots,\boldsymbol{\theta}_{M}]$ represent the set of trainable parameters for the ansatz circuit containing $M$ layers. Each layer is composed of unitary operations that depend on the parameters $\boldsymbol{\theta}_{a}$. Each layer parameter set $\boldsymbol{\theta}_{a\in\{1,\cdots,M\}}$ includes two subsets of trainable parameters, $\boldsymbol{\theta}_{a_{1}}\in\mathbb{R}^{2\times m}$ and $\boldsymbol{\theta}_{a_{2}}\in\mathbb{R}^{m-1}$, where $m$ is the number of qubits.
@@
$\text{$RX_{q}$}(\theta_{a_{1}}^{(q_{1})})=\exp\left(-i\frac{\theta_{a_{1}}^{(q_{1})}}{2}X_{q}\right),$ (15)
$\text{$RZ_{q}$}(\theta_{a_{1}}^{(q_{2})})=\exp\left(-i\frac{\theta_{a_{1}}^{(q_{2})}}{2}Z_{q}\right)$ (16)
where $\theta_{a_{1}}^{(q_{1})}$ and $\theta_{a_{1}}^{(q_{2})}$ are the two parameters from $\boldsymbol{\mathbf{\theta}_{a_{1}}}$ for the qubit $q$.
@@
$\text{CRZ}_{q,q+1}(\phi_{q,q+1})=\begin{pmatrix}1&0&0&0\\ 0&1&0&0\\ 0&0&e^{-i\phi_{q,q+1}/2}&0\\ 0&0&0&e^{i\phi_{q,q+1}/2}\end{pmatrix}$ (17)
@@
$\lvert\psi_{\text{ansatz}}\rangle=\prod_{q=1}^{m-1}\text{$CRZ_{q}$}(\phi_{q,q+1})\prod_{q=1}^{m}\text{$RZ_{q}$}(\theta_{a_{1}}^{(q_{2})})\text{$RX_{q}$}(\theta_{a_{1}}^{(q_{1})})\lvert\psi_{\text{enc}}\rangle$ (18)
@@ III-D Complexity Analysis of the Model
The number of qubits, $m$, required by the model is governed by the size of the vocabulary $|V|$. If each word in the vocabulary is represented by a computational basis state, the number of qubits is proportional to the logarithm of the vocabulary size,
@@
  • Ansatz Block Complexity: The ansatz block is repeated $n_{\text{repeats}}$ times. For a system of $m$ qubits, each layer of the ansatz consists of $2m$ single-qubit rotation gates and m1m-1 controlled entangling gates (such as CNOT or CRZ). Thus, the number of gates in a single ansatz layer is approximately $G_{\text{layer}}=3m-1$. The total number of gates required for the ansatz block with $n_{\text{repeats}}$ repetitions is
@@
$G_{\text{total}}=G_{\text{ansatz}}+G_{\text{context}}=(3m-1)(M+L)+m.$ (23)
IV Computing the Context Matrices and Vectors
This section details approaches to generate unique context matrices. These methods are grounded in various transformations, such as the hash function [61], exponential decay, and sinusoidal functions[62, 63]. The aim of these functions is to encode word relationships within context windows, producing matrices that capture word identities, distance and positional information, and semantic-syntactic features. Then, these matrices are reshaped to $\mathbf{\tilde{C}}\in\mathbb{R}^{2m\times\lceil{n^{2}/2m}\rceil}$ and padded with zeros to match the layers.
IV-A Exponential Decay with Sinusoidal Encoding
This method defines the context matrix $\mathbf{C}\in\mathbb{R}^{n\times n}$ for a context window of $n$ words by incorporating both positional distance and semantic relationships through sinusoidal transformations. For words at positions $i$ and $j$ in the context window, each element $c_{ij}$ is computed as
@@ IV-A Exponential Decay with Sinusoidal Encoding
where:
  • $\theta_{i}=\text{idx}_{i}\cdot\frac{2\pi}{|V|}$ is the normalized angle for word at position $i$.
  • $\text{idx}_{i}$ is the vocabulary index of the word at position $i$.
  • |ij||i-j| represents the positional distance between words.
  • $\alpha>0$ controls the exponential decay rate with distance.
  • $\omega>0$ determines the frequency of sinusoidal components.
  • $|V|$ is the vocabulary size or maximum index value.
This formulation combines three key components: (1) an exponential decay term that captures the positional proximity between words, (2) a sinusoidal encoding that provides distinct representations for different words, and (3) an index-based uniqueness term that ensures different words receive distinct representations even when they appear at the same relative positions.
IV-B Index-Based Diagonal Encoding
The second approach introduces a specialized treatment for diagonal elements to emphasize self-relationships while maintaining positional decay for off-diagonal elements. The context matrix $\mathbf{C}\in\mathbb{R}^{n\times n}$ is defined as
where:
  • $\theta_{i}=\text{idx}_{i}\cdot\frac{2\pi}{|V|}$ is the normalized angle
  • $\log(1+\text{idx}_{i})$ provides a logarithmically scaled value for diagonal element and assuming that indexes start from 0.
  • Other parameters follow the definitions from the previous method.
The logarithmic scaling of diagonal elements ensures that words with higher indices (presumably unknown or less common words) receive stronger self-relationships, which can be valuable for preserving the significance of rare terms in the context representation.
IV-C Positional Phase Shift Encoding
This method introduces position-dependent phase shifts in the sinusoidal components, creating a rich representation that encodes both absolute positions and relative distances. The context matrix $\mathbf{C}\in\mathbb{R}^{n\times n}$ is defined as
$c_{ij}=e^{-\alpha|i-j|}\sin(\omega i+\delta\theta_{i})+\theta_{i}$ (26)
where:
  • $\delta>0$ is a phase shift scaling parameter.
  • $\omega i$ introduces a position-dependent base phase.
  • $\delta\theta_{i}$ adds a word-specific phase shift.
  • Other parameters follow previous definitions.
This formulation effectively creates a unique encoding for each word based on both its identity and its position in the context window, enabling the model to distinguish between identical words appearing at different positions.
IV-D Hash-Based Modulation Encoding
To enhance the uniqueness of representations and reduce potential collisions between words with similar indices, this method introduces a hash-based modulation approach. The context matrix $\mathbf{C}\in\mathbb{R}^{n\times n}$ is defined as
$c_{ij}=e^{-\alpha|i-j|}\sin(\omega i+h_{i})+\theta_{i}$ (27)
where:
  • $h_{i}=(\text{idx}_{i}\cdot p)\bmod N$ is a simple hash function
  • $p$ is a prime number
  • NN is the hash space size
  • Other parameters follow previous definitions
The hash-based modulation introduces an additional layer of uniqueness that is particularly valuable for handling large vocabularies where index-based collisions might otherwise occur. The use of prime numbers in the hash function helps ensure a good distribution of hash values across the vocabulary.
IV-E Positional Angular Shift Vector Representation
While the previous methods focused on matrix representations, a context-to-vector method introduces a vector-based context representation that efficiently captures word relationships. A context window of $n$ words, is defined as context vector $\mathbf{v}\in\mathbb{R}^{n}$ where each element $v_{i}$ is
$v_{i}=\omega\theta_{i}$ (28)
where:
  • $\theta_{i}=\text{idx}_{i}\cdot\frac{2\pi}{|V|}$ is the normalized angle
  • $\omega>0$ determines the frequency of angular encoding
  • $\text{idx}_{i}$ is the vocabulary index of the word at position $i$
This vector representation provides a more compact and straightforward encoding that emphasizes word identities through angular transformations while incorporating positional information implicitly through the qubits ordering or indices.
Each of these methods offers different trade-offs between computational complexity, representational capacity, and numerical stability.
V Evaluation of QCSE
The QCSE model was implemented and tested using PennyLane [^1] https://pennylane.ai/ in two separate evaluations. The first (Section V-A) assessed five context encoding methods (Section IV) to identify the most effective approach. The second (Section V-B) compared the quantum embedding model with word2vec’s CBOW [64] to examine potential quantum advantages in QCSE.
V-A Evaluation of Context Encoding Methods
This section outlines the evaluation setup and results for assessing the QCSE model’s performance with different context encoding methods across two corpora.
V-A1 Evaluation Setup
The evaluations were structured as follows. Two datasets were used:
  • English: a vocabulary size of $31$ and 656 center-context word pairs.
  • Fulani (a low-resource African language) [65]: a vocabulary size of $27$ and 128 word-context pairs.
The model used 5 qubits, with each word encoded as a quantum state. The quantum ansatz had 3 layers (14 trainable parameters each, totaling 42 parameters). A context window of 2 (4-word span) was applied, following the method in Figure 1.
V-A2 Results and Discussion
(a) English dataset (656 pairs)
(b) Fulani dataset (128 pairs)
Figure 4: Training loss across methods: (a) Exponential Decay Sinusoidal, (b) Index-based Diagonal, (c) Positional Phase Shift, (d) Hash-based, (e) Positional Angular Shift. Fulani represents a low-resource African language.
The training process optimizes embedding similarity via cross-entropy loss minimization, and we compare convergence rates to find the method with the fastest stable loss reduction. The results, Figure 4(a) and Figure 4(b) for English and Fulani datasets respectively, demonstrate that Exponential Decay Sinusoidal and the Positional Phase Shift encoding methods consistently yield the best performance across both datasets. In the English dataset, the Hash-based and Index-based Diagonal methods achieve the lowest final training loss, around $4.87$. The Positional Angular Shift method performs worst, fluctuating around $5.2$, indicating poorer learning of context. These trends suggest that smooth, structured encodings—such as sinusoidal or phase-based—enable more effective learning of contextual relationships.
In the Fulani dataset (Figure 4(b)), where only 128 word-context pairs were used. The Exponential Decay Sinusoidal encoding reaches a final loss of about $6.1$, the lowest among all methods, while the Hash-based method stabilizes around $6.2$. Other methods like Index-based Diagonal and Hash-based settle around $6.3$, while Positional Angular Shift remains the worst performer at approximately $6.6$ throughout training. The sharper performance gap in the low-resource setting highlights the robustness of the Exponential Decay Sinusoidal encoding, as it can still effectively represent positional context even with limited data. This could support its suitability for both high- and low-resource NLP applications within quantum embedding models.
The success of these methods can be attributed to their inherent compatibility with quantum systems. Exponential Decay Sinusoidal leverages periodicity and decay to encode positional information. This performance underscores the importance of encoding methods that balance quantum native operations with classical NLP principles (e.g., positional awareness), especially when scaling across dataset sizes. The next section analyses the performance of the QCSE model using the exponential decay sinusoidal method for context encoding.
V-B Analysis of QCSE model performance
This section presents a comprehensive analysis of QCSE model performance compared to classical word embedding approaches, specifically the well-known CBOW model[64]. The evaluation focuses on parameter efficiency, scalability characteristics.
V-B1 Evaluation Setup
Evaluations were carried out on a data set of vocabulary size 34, comprising 300 sentences (1200 words total), from which 1200 center-context word pairs were extracted. The data set was systematically partitioned into training and testing subsets to ensure robust evaluation. The QCSE model architecture employs 6 qubits with varying layer depths (1-8 layers), while the classical CBOW baseline utilizes vector dimensions of 20 and 50. All models were evaluated using similarity metrics on word relationship tasks, providing a standardized measure of semantic representation quality.
V-B2 Results and Discussion
The dataset is split into training and test sets. The model is trained to maximize similarity between true and predicted embeddings via cross-entropy loss. During validation, accuracy is measured using the hamming distance to compute the similarity scores by checking if the predicted word or embedding matches the center of the context. The predictions were considered accurate if the similarity between quantum embeddings exceeded $0.90$, equivalent to a Hamming distance of at most 1 qubit difference. Table I shows that the QCSE model exhibits distinctive performance characteristics as a function of the depth of the circuit. In this evaluation, the optimal architecture is identified with 6 layers, achieving 37.82% accuracy with only 102 trainable parameters and outperforms the classical model (37.82% vs 34.78%). This represents a clear performance peak, with both shallower and deeper architectures showing reduced effectiveness, as reflected in the curves. The performance curve reveals several critical insights into quantum word embedding behavior: (1) Shallow Circuit Limitations: Single-layer QCSE (9.58% accuracy) demonstrates insufficient expressibility for capturing complex semantic relationships. The rapid improvement from 1 to 4 layers (9.58% to 36.67%) indicates that quantum circuits require adequate depth to develop meaningful word representations. (2) Optimal Depth Identification: The performance plateau between 4-6 layers, with peak performance at 6 layers (37.82%), suggests an optimal balance between circuit expressibility and trainability. This finding is crucial for practical quantum NLP applications, as it identifies the minimal circuit depth required for competitive performance with respect to the data set size. (3) Overparameterization Effects: Performance degradation beyond 6 layers (37.82% to 34.58% at 8 layers) indicates quantum circuit over-parameterization, analogous to classical deep learning phenomena. This suggests that quantum models are subject to similar optimization challenges as their classical counterparts, requiring careful architectural design. The results of the classical CBOW model reveal interesting scalability characteristics that further emphasize the quantum parameter efficiency. Counter-intuitively, increasing the CBOW vector dimension from 20 to 50 (increasing parameters from 680 to 1700) with the same epochs resulted in decreased performance from 34.78% to 20.00%. This performance degradation suggests that, just like the quantum model, classical models suffer from optimization difficulties when scaled inappropriately for limited training data[64].
These results demonstrate the parameter utilization efficiency of the QCSE model compared to the CBOW model. The optimal QCSE model (6 layers, 102 parameters) outperforms the classical CBOW model with 680 parameters, achieving 3.04% higher accuracy with 85% fewer parameters. This represents a parameter efficiency improvement when measured by accuracy per parameter. When comparing against the larger CBOW configuration (50-dimensional, 1700 parameters), the quantum model achieves 17.82% higher accuracy (37.82% vs. 20.00%) while using 94% fewer parameters. This contrast highlights the ability of the quantum model to capture semantic relationships with fewer parameters. The parameter efficiency becomes particularly relevant in resource-constrained environments where parameter count directly impacts computational overhead and memory requirements.
TABLE I: Comparative Performance Analysis of Pretrained QCSE and Classical Continuous Bag of Words (CBOW) Models
Model Number of Number of Number of Number of Accuracy
layers qubits/Vec trainable Training [%]
tor dimension params words (similarity
score)
QCSE 1 6 17 1200 9.58
2 6 34 1200 27.09
3 6 51 1200 32.50
4 6 68 1200 36.67
5 6 85 1200 32.50
6 6 102 1200 37.82
7 6 119 1200 36.67
8 6 136 1200 34.58
CBOW 20 680 1200 34.78
50 1700 1200 20.00
The training loss curves presented in Figure 5 provide insights into the learning dynamics, convergence behavior, and optimization characteristics of both quantum and classical word embedding models. The QCSE models exhibit different training behaviors as a function of circuit depth, with each layer configuration demonstrating unique convergence characteristics that correlate with final performance outcomes.
The single-layer quantum circuit displays unstable training dynamics with oscillatory behavior throughout the training process. The loss remains constant around 8.8, never achieving convergence. This behavior directly corresponds to the poor final accuracy of 9.58%, indicating that single-layer quantum circuits lack sufficient expressibility to capture meaningful semantic relationships. As circuit depth increases from 2 to 4 layers, we observe a progressive improvement in training stability and convergence behavior. The 2-layer circuit shows rapid initial convergence from approximately 6.4 to 6.17 within the first 10 epochs, followed by stable plateauing around 6.2. This convergence pattern correlates with the performance improvement to 27.09% accuracy. The 3-layer configuration demonstrates stable training dynamics, with a loss reduction from approximately 6.0 to 5.55 and achieving 32.5% accuracy. The 4-layer model exhibits stable training trajectory among shallow circuits, with consistent loss reduction from about 5.9 to 5.55 and minimal oscillations, corresponding to 36.67% accuracy.
The 6-layer quantum circuit, which achieved the highest accuracy of 37.82%, displays a best training dynamics characterized by monotonic loss reduction from approximately 6.0 to 5.5. The convergence curve shows three distinct phases: rapid initial descent (epochs 0-20). Deeper quantum circuits (7-8 layers) exhibit similar final loss values but with slightly reduced accuracy.
The 20-dimensional CBOW model (680 parameters and 34.78% accuracy) exhibits a training dynamics with fluctuations throughout the training process. The loss fluctuates between approximately 6.9 and 6.35, showing periodic variations. The higher-dimensional CBOW model (1700 parameters) shows similar training oscillations but achieves significantly lower performance (20.00% accuracy). The training curve displays persistent oscillations with loss values fluctuating around 6.4, indicating that increased model dimensionality does not translate to improved optimization or performance for this dataset size[64].
Compared to classical models such as word2vec[64], GloVe[44, 32], which typically require high-dimensional vectors and large data sets, the proposed quantum model demonstrates the ability to perform well on large and small datasets. Furthermore, compared to other quantum embedding models that rely on classical embeddings as input[47, 32], QCSE model learns directly from scratch. This significantly reduces the dependence on classical pre-trained embeddings and highlights the efficiency of using quantum language modeling.
This suggests that quantum parameterized circuits, when properly designed, may possess more favorable optimization geometries than their classical counterparts. Our quantum model can achieve stable convergence with significantly fewer parameters. The proposed quantum circuits require fewer epochs to reach stable convergence, with most QCSE models achieving plateau performance within 10-30 epochs compared to the continued instability of classical models throughout 50 epochs.
Figure 5: Training loss curves that analyzes both QCSE and CBOW behaviors.

### v2

@@ III-B2 Ansatz Circuit and Trainable Parameters
The second component of the QCSE model is the ansatz circuit, which comprises trainable parameters. The architecture of the ansatz circuit is shown in Figure 3 in the red block. Let $\mathbf{\Theta}=[\boldsymbol{\theta}_{1},\boldsymbol{\theta}_{2},\dots,\boldsymbol{\theta}_{M}]$ represent the set of trainable parameters for the ansatz circuit containing $M$ layers. Each layer is composed of unitary operations that depend on the parameters $\boldsymbol{\theta}_{a}$. Each layer parameter set $\boldsymbol{\theta}_{a\in\{1,\cdots,\text{ }M\}}$ includes two subsets of trainable parameters, $\boldsymbol{\theta}_{a_{1}}\in\mathbb{R}^{2\times m}$ and $\boldsymbol{\theta}_{a_{2}}\in\mathbb{R}^{m-1}$, where $m$ is the number of qubits.
@@
$\text{$RX_{q}$}(\theta_{a_{1}}^{(q,1)})=\exp\left(-\text{i}\frac{\theta_{a_{1}}^{(q,1)}}{2}X_{q}\right),$ (15)
$\text{$RZ_{q}$}(\theta_{a_{1}}^{(q,2)})=\exp\left(-\text{i}\frac{\theta_{a_{1}}^{(q,2)}}{2}Z_{q}\right)$ (16)
where $\theta_{a_{1}}^{(q,1)}$ and $\theta_{a_{1}}^{(q,2)}$ are the two parameters from $\boldsymbol{\mathbf{\theta}_{a_{1}}}$ for the qubit $q$.
@@
$\text{CRZ}_{q,q+1}(\phi_{q,q+1})=\begin{pmatrix}1&0&0&0\\ 0&1&0&0\\ 0&0&e^{-\text{i}\phi_{q,q+1}/2}&0\\ 0&0&0&e^{\text{i}\phi_{q,q+1}/2}\end{pmatrix}$ (17)
@@
$\begin{split}\lvert\psi_{\text{ansatz}}\rangle=&\text{CRZ}_{m-1}(\phi_{m-1,m})\cdots\text{CRZ}_{1}(\phi_{1,2})\\ &\cdot\left(\bigotimes_{q=1}^{m}\text{RZ}_{q}(\theta_{a_{1}}^{(q,2)})\text{RX}_{q}(\theta_{a_{1}}^{(q,1)})\right)\lvert\psi_{\text{ent}}\rangle\end{split}$ (18)
@@ III-D Complexity Analysis of the Model
The number of qubits $m$ required by the model is governed by the size of the vocabulary $|V|$. If each word in the vocabulary is represented by a computational basis state, the number of qubits is proportional to the logarithm of the vocabulary size,
@@
  • Ansatz Block Complexity: The ansatz block is repeated $n_{\text{repeats}}$ times. For a system of $m$ qubits, each layer of the ansatz consists of $2m$ single-qubit rotation gates and m1m-1 controlled entangling gates (such as CNOT or CRZ). Thus, the number of gates in a single ansatz layer is $G_{\text{layer}}=3m-1$. The total number of gates required for the ansatz block with $n_{\text{repeats}}$ repetitions is
@@
$G_{\text{total}}=G_{\text{ansatz}}+G_{\text{context}}=(3m-1)(M+L).$ (23)
IV Context Encoding: Exponential Decay with Sinusoidal Method
The choice of context representation is fundamental to the quantum embedding framework. Unlike classical word embedding models that typically employ flat vector representations for context words, we propose encoding the context as a structured matrix to capture richer relational information between words within the context window. This design is motivated by several considerations: (1) matrix representations naturally encode pairwise relationships, enabling the model to capture not just individual word identities but also their interactions; (2) the structured format provides multiple dimensions for encoding positional, semantic, and distance information simultaneously; (3) this approach draws inspiration from positional encoding mechanisms in transformer architectures [vaswani2017attention], which have proven effective at capturing word order and relative positions.
Among the various encoding strategies we explored, the exponential decay with sinusoidal encoding method demonstrated consistent and robust performance across different model depths. This method defines the context matrix $\mathbf{C}\in\mathbb{R}^{n\times n}$ for a context window of $n$ words, where each element $c_{ij}$ capturing the relationship between words at positions $i$ and $j$ is computed as:
@@ IV Context Encoding: Exponential Decay with Sinusoidal Method
where $\theta_{i}=\text{idx}_{i}\cdot\frac{2\pi}{|V|}$ represents the normalized angle for the word at position $i$, with $\text{idx}_{i}$ denoting its vocabulary index and $|V|$ the vocabulary size. The parameters $\alpha>0$ and $\omega>0$ control the exponential decay rate based on the word index, distance between words, and the changing position of words.
This formulation integrates three complementary encoding mechanisms. First, the exponential decay term $e^{-\alpha|i-j|}$ captures positional proximity: words closer in the context window receive stronger weighting, reflecting the linguistic intuition that nearby words typically have strong relationships. Second, the sinusoidal encoding $\sin(\omega\theta_{i})\cos(\omega\theta_{j})$ provides distinct representations for different vocabulary items through periodic functions, similar to positional encodings in transformers [vaswani2017attention, shaw2018self]. Third, the additive term $\theta_{i}$ ensures that each word maintains a unique signature based on its vocabulary position, preventing representational collapse even when words appear at identical relative positions in different contexts.
After construction, the context matrix is reshaped to $\mathbf{\tilde{C}}\in\mathbb{R}^{2m\times\lceil n^{2}/2m\rceil}$ and padded with zeros as needed to match the circuit architecture, where $m$ denotes the number of qubits. This reshaping distributes the context information across multiple encoding layers, allowing the quantum circuit to process the full relational structure through sequential gate operations.
While this exponential decay method proved most effective in our experiments, the design space for context encoding remains explorable. We investigated four alternative encoding strategies—index-based diagonal encoding, positional phase shift encoding, hash-based modulation, and positional angular shift—each offering different trade-offs in terms of representational capacity, computational efficiency, and numerical properties. These alternative methods and their comparative performance are detailed in the Appendix.
V Evaluation and Results
V-A Setup
We evaluated the proposed QCSE model against classical word embedding baselines using two datasets: an English corpus (110 sentences, 550 tokens, 80-word vocabulary, requiring 7 qubits) and a Fulani[10.11648] corpus (20 sentences, 120 tokens, 26-word vocabulary, requiring 5 qubits). The primary objective is to assess whether the quantum model can effectively capture contextual relationships between center–context word pairs under limited model capacity.
Due to differences in output encoding schemes and their corresponding loss functions, distinct hyperparameter configurations were employed for quantum and classical models. Quantum models used a learning rate of 0.0003 and L2 regularization ($\lambda=0.001$), minimizing binary cross-entropy loss to accommodate bitwise output encoding. Classical baselines (CBOW and Skip-gram variants) employed categorical cross-entropy loss with learning rate 0.03 and L2 regularization ($\lambda=0.0001$), reflecting their probabilistic multi-class formulation. All experiments used an 80-20 train-test split, a context window size of 4, and were trained for 50 epochs.
The accuracy was calculated as follows. For quantum models, predicted probabilities were first converted into binary bitstrings using a threshold of 0.5. The similarity score between predicted and target bitstrings was then calculated as the proportion of matching bits, and a prediction was considered correct if this similarity score was at least 0.5. In the same way, for classical models, the accuracy was defined as the percentage of context words whose predicted representations achieved a similarity score of at least 0.5 with the target representations.
V-B Discussion
Table I presents the accuracy results for both quantum and classical models on the English dataset. The quantum models, implemented using the exponential decay sinusoidal encoding method, scale linearly with depth at 21 parameters per layer (ranging from 21 to 168 parameters for 1-8 layers). Classical baselines range from 8,000 to 48,000 parameters depending on embedding dimension.
TABLE I: Accuracy comparison on English dataset—QCSE (exponential decay method) vs. classical models
Model Parameters Accuracy (%)
Classical Models
CBOW-50 8,000 65.5
CBOW-80 12,800 70.0
CBOW-100 16,000 68.1
CBOW-200 32,000 64.6
CBOW-300 48,000 63.7
SkipGram-50 8,000 53.8
SkipGram-80 12,800 62.1
SkipGram-100 16,000 58.8
SkipGram-200 32,000 58.6
SkipGram-300 48,000 58.2
Quantum Models (Exp. Decay Sinusoidal)
QCSE-1L 21 50.4
QCSE-2L 42 64.6
QCSE-3L 63 67.3
QCSE-4L 84 70.8
QCSE-5L 105 67.3
QCSE-6L 126 66.4
QCSE-7L 147 84.1
QCSE-8L 168 85.8
Table II shows results on the smaller Fulani dataset, which provides insight into model behavior under more limited data conditions.
TABLE II: Accuracy comparison on Fulani dataset—QCSE (exponential decay method) vs. classical models
Model Parameters Accuracy (%)
Classical Models
CBOW-50 2,600 76.9
CBOW-80 4,160 61.5
CBOW-100 5,200 57.7
CBOW-200 10,400 53.8
CBOW-300 15,600 61.5
SkipGram-50 2,600 47.1
SkipGram-80 4,160 46.5
SkipGram-100 5,200 51.3
SkipGram-200 10,400 55.1
SkipGram-300 15,600 54.2
Quantum Models (Exp. Decay Sinusoidal)
QCSE-1L 15 46.2
QCSE-2L 30 34.7
QCSE-3L 45 46.2
QCSE-4L 60 69.2
QCSE-5L 75 77.0
QCSE-6L 90 73.1
QCSE-7L 105 80.8
QCSE-8L 120 61.5
The quantum model using exponential decay sinusoidal encoding method demonstrates a clear progression in model capability with increasing depth. On the English dataset (Table I), single-layer models achieve only 50.4% accuracy indicating insufficient capacity to capture meaningful contextual relationships. Performance improves at 2-4 layers (64.6-70.8%), with the 4-layer configuration matching the best classical baseline (CBOW-80 at 70.0%) while using only 84 parameters compared to 12,800. In deeper configurations: the 7-layer model achieves 84.1% accuracy, and the 8-layer model reaches 85.8%, representing a 15.8 percentage point improvement over the best classical baseline while using 76× fewer parameters (168 vs. 12,800). This improvement at 7-8 layers suggests that the quantum model benefits from increased expressivity when sufficient circuit depth is available to exploit the relational structure encoded in the context matrices. However, the 5-layer (67.3%) and 6-layer (66.4%) models perform slightly worse than the 4-layer configuration. There may be several reasons for this phenomenon. One possibility is that this intermediate fluctuation may reflect an optimization instability. As circuit depth increases, the optimization landscape of some layers becomes more complex, which can lead to suboptimal local minima. Performance recovers at greater depths, suggesting that once the model surpasses the transitional region, the capacity becomes effectively utilized. Another possibility is that this phenomenon could be implicitly influenced by the dual thresholding mechanisms in the accuracy computation, in the same way that this inconsistency was observed with loss and accuracy. First, predicted probabilities are converted to bits using a 0.5 threshold; a probability shifting from 0.50 to 0.49 across epochs changes the bit assignment despite maintained model confidence, affecting similarity scores. Second, predictions are counted as correct only when the final similarity score exceeds 0.5. These thresholds can cause temporary accuracy decreases even as the model continues learning, as evidenced by subsequent accuracy recovery in the training curves.
On the Fulani dataset (Table II), the optimal depth shifts to 7 layers (80.8%), with the 8-layer model exhibiting degraded performance (61.5%). This suggests dataset-dependent capacity requirements: smaller datasets may not provide sufficient training information to effectively optimize deep circuits, leading to capacity saturation or optimization difficulties beyond a certain model depth. In Figure 4, for the English corpus, performance increases to a maximum at 168 parameters (85.8%) and for the smaller Fulani corpus, the peak occurs earlier at 105 parameters (80.8%), followed by a decline. This shift in optimal parameter count (168 → 105) reflects adaptation to dataset size. Quantum models remain parameter-efficient, but require alignment between circuit depth and dataset size to achieve optimal performance. This degradation resembles the capacity–generalization trade-off observed in classical models [mikolov2013efficient]. Classical models exhibit a known phenomenon in classical embedding  [mikolov2013efficient]: when all other parameters remain constant, increasing the embedding dimension tends to degrade performance. On the English dataset, CBOW achieves comparable accuracy at 80 dimensions (70.0%), with performance declining at higher dimensions: 68.1% (100-dim), 64.6% (200-dim), and 63.7% (300-dim). Skip-gram shows a similar pattern, with performance dropping from 62.1% (80-dim) to 58.2% (300-dim). This behavior reflects a fundamental challenge in classical embedding models: higher-dimensional representations increase model capacity, but without sufficient training data, they become increasingly difficult to optimize and prone to overfitting. On the moderate-sized English corpus (550 tokens), models with 200-300 dimensional embeddings have too many parameters (32,000-48,000) to effectively constrain through the available training samples, resulting in poor generalization despite lower training loss. The Fulani dataset follows the same effect: CBOW-50 achieves 76.9% accuracy with 2,600 parameters, but performance collapses at higher dimensions (53.8% for 200-dim, 61.5% for 300-dim). This demonstrates that classical embedding models, designed and optimized for large-scale corpora with millions of training examples, struggle to adapt their capacity to small-dataset regimes. The quantum model exhibits similar challenges. For example, on the Fulani dataset, the 8-layer quantum model drops to 61.5% despite the 7-layer model achieving 80.8%. This suggests that with fewer parameters (120 vs. 15,600 for CBOW-300), quantum circuits could still exceed optimal capacity for very small datasets.
Figure 4: Accuracy versus trainable parameter count in QCSE using Exponential Decay Sinusoidal Method on English and Fulani datasets.
Analysis of training dynamics indicates that deeper circuits (7 to 8 layers) start with higher initial loss but exhibit slightly sustained descent, achieving higher accuracy. Some shallow models converge quickly with lower final loss, but yield poor accuracy. The corresponding analysis of loss trajectories, convergence regimes, and comparisons together with all alternative encoding methods is provided in the Appendix A.